### v1

# arXiv:2603.01466v1[math.FA] 02 Mar 2026

Abstract.
Differently from the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework often reveals information about the algebraic structure. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse. Our results not only utilize the violation of bilocal inequalities to reveal the structural properties of von Neumann algebras, but can also be applied to quantum mechanics and quantum field theory.
Corresponding author
Key words and phrases: von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
1. Introduction
Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
Moreover, Bell nonlocality in quantum field theory (QFT) serves as a crucial bridge connecting quantum information science with fundamental physics [16, 17]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [19, 18]. Furthermore, the tension between nonlocal quantum phenomena and the locality assumptions of general relativity offers profound insights for quantum gravity research, suggesting that spacetime itself may be an emergent property rather than a fundamental aspect of reality [20, 21, 22].
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [23, 24, 25]. This differs from the traditional quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [27, 26, 28]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model. In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, denoted as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ However, it should be pointed out that relying solely on the above tensor-product model to discuss quantum information problems has two drawbacks [29, 30, 31]: 1) It fails to provide a universal framework for accurately describing phenomena in infinite-dimensional systems. 2) It is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, where $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$, and $\mathcal{M}_{A},\mathcal{M}_{B}$ are two mutually commuting von Neumann subalgebras of $\mathcal{M}$. The notation $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$. However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{A}\otimes\mathcal{M}_{B}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [33, 32, 34]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [34] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product model algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [34, 35, 36, 37].
Overall, Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [17]. Summers et al. introduces the maximal Bell correlation and proves that its value is bounded by $\sqrt{2}$ in mutually-commuting von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [38]. Translating these bounds into the vacuum representation of algebraic quantum field theory, he shows that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges and decay exponentially with mass-governed distance [39, 40, 41, 42, 43]. These works expose a new algebraic invariant¡ªthe Bell correlation invariant¡ªthat distinguishes infinitely many isomorphism classes of mutually-commuting von Neumann algebra pairs and links maximal violation to the presence of the hyperfinite type-${\rm II}_{1}$ factor [17].
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [44, 45, 46, 47, 48, 49]. To date, Bell-type inequalities have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [44, 50], chain configurations [51, 52], star topologies [53, 54], polygon structures [55, 56, 57], tree-shaped networks [58, 59, 60], arbitrary acyclic networks [46, 47, 61], and arbitrary $k$-independent networks [62]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [63, 64]. Furthermore, the interplay between graph theory and quantum correlations has attracted considerable attention [65, 66]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum topologies [67]. Interestingly, the notion of bi-locality in the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al [68], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [69].
In this paper we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Figure 1. An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. All the three parties can perform measurements (or inputs) on their system, labeled $x$, $y$ and $z$ for Alice, Bob and Charles. After measuring the three parties, they obtain outcomes denoted $a$, $b$ and $c$, respectively. Here, $x,y,z,a,b,c\in\{0,1\}$. Bob’s measurement $y$ might correspond to a joint measurement on the two systems that he receives from each source. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, the joint probability distribution $p(abc|xyz)$ is bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu q_{1}(\lambda)q_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ ($\mu$) characterizes the joint hidden variable of the system produced by the source $\rho_{AB}$ ($\rho_{BC}$) [70, 71].
Let $A_{x},B_{y},C_{z}$ denote the measurements of Alice, Bob, and Charles with inputs $x$, $y$, $z$, respectively. Note that the middle party Bob always performs joint measurement. Then, the bilocal inequality can be written as:
(2.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
where
$I\equiv\sum_{x,z}\langle A_{x}B_{0}C_{z}\rangle=\langle(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\rangle,$
$J\equiv\sum_{x,z}(-1)^{x+z}\langle A_{x}B_{1}C_{z}\rangle=\langle(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\rangle$
and $\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz)$ [72].
In addition, one can choose different observables to construct different bilocal inequalities. For instance, the violation of the inequality
(2.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [73]. Of course, different inequalities exhibit varying capabilities in detecting nonlocality.
Mutually-commuting von Neumann algbera models. In QFT, the observables for Alice, Bob and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{A},\ \mathcal{M}_{B},\ \mathcal{M}_{C}$. In general, they are all required to be type III von Neumann algebra factors. Recall two von Neumann algebras satisfy Haag duality means that each algebra equals the commutant of the other. Haag duality encode quantum nonlocality intrinsically within their infinite-dimensional non-abelian structure through their type III factor property, which enforces that local algebras of spacelike separated regions are entangled with trivial relative commutants [74, 75]. Here, we do not restrict von Neumann algebras to III type factors, and generalize the Haag duality condition to mutual commutativity, proposing a von Neumann algebra model with mutual commutativity. Therefore, our model covers both scenarios of non-relativistic quantum mechanics and quantum field theory. The idea of this model is similar to that in Refs. [69, 68].
Definition 1. (Ternary mutually-commuting von Neumann algebra models) Let $\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C}$ be von Neumann subalgebras of $\mathcal{B}(H)$ over some Hilbert space $H$, and they are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{A,B,C\}$, where $\mathcal{M}_{j}^{\prime}$ is the communicator of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M_{ABC}}=(\mathcal{M_{A}}\vee\mathcal{M_{B}}\vee\mathcal{M_{C}})^{\prime\prime}.$
We refer to the above model as the MCvNA model. When $\mathcal{M}_{ABC}=\mathcal{M}_{A}\otimes\mathcal{M}_{B}\otimes\mathcal{M}_{C}$, it is called the tensor product algebra model. Here, in order to correspond to the Hermitian property of quantum observables, these three algebras are always restricted to consist of self-adjoint elements.
We point out that the von Neumann algebra model defined above is intended for the study of algebraic structural properties; therefore, we have adopted the most general definition, which encompasses but does not correspond to models of quantum mechanics or quantum field theory.
We intend to use the above model of three mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between parties $A$ and $C$ in the network. Mathematically, this independence can be described by the following formula. Let $\tau$ be a state in the dual space $\mathcal{M}_{ABC}^{*}$. In correspondence with the bilocal scenario, we always require that
() $\tau(ac)=\tau(a)\tau(c)$
for any $a\in\mathcal{M}_{A},c\in\mathcal{M}_{C}$. We call it as the independent condition (*). This assumption will be used throughout this paper.
3. Bilocal inequalities and their bounds
Following the assumptions in Section 2, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ is defined as
$p(\alpha\beta\gamma|xyz)=\tau({A}_{\alpha|x}{B}_{\beta|y}{C}_{\gamma|z}),$
where the measurement operators ${A}_{\alpha|x}\in\mathcal{M}_{A},\ {B}_{\beta|y}\in\mathcal{M}_{B},\ {C}_{\gamma|z}\in\mathcal{M}_{C}$, $\{x,y,z\}$ and $\{\alpha,\beta,\gamma\}$ denote the inputs and outputs of measurement operators with $x,y,z,\alpha,\beta,\gamma\in\{0,1\}$. Let $a_{x}\in\mathcal{M}_{A}$, $b_{y}\in\mathcal{M}_{B}$, $c_{z}\in\mathcal{M}_{C}$ be operators such that $I\leq a_{x},b_{y},c_{z}\leq I$ for any $x,y,z\in\{0,1\}$, and $a_{x}=\sum_{\alpha=0}^{1}(-1)^{\alpha}A_{\alpha|x},b_{y}=\sum_{\beta=0}^{1}(-1)^{\beta}B_{\beta|y},c_{z}=\sum_{\gamma=0}^{1}(-1)^{\gamma}C_{\gamma|z}$. Then denote $I_{\tau}=\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1})),\ J_{\tau}=\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))$. In the MCvNA model, analogous to Ineq. (2.1), we define
(3.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
Analogous to Ineq. (2.2), under the same assumption, we have the following
(3.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|.$
where, $K_{\tau}=\tau(a_{0}(b_{0}+b_{1})c_{0}),\ L_{\tau}=\tau(a_{1}(b_{0}-b_{1})c_{1})$.
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively. We mention that all theorems in this section are based on the independence assumption (*).
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the maximum upper bound that the quantity (3.1) can achieve is $2\sqrt{2}$. This is analogous to the case in non-relativistic quantum mechanics, where the maximal violation of the bilocal value based on quantum resources is $2\sqrt{2}$.
Theorem 1. In the MCvNA model, $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{ABC}\rightarrow H_{\tau}$ and a cyclic vector $\Omega\in H_{\tau}$ such that the set $\{\pi_{\tau}(d)\Omega:\ d\in\mathcal{M}_{ABC}\}$ is dense in $H_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
$=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$\leq\sqrt{2}\sqrt{|\tau(b_{0}(a_{0}+a_{1})(c_{0}+c_{1}))|+|\tau(b_{1}(a_{0}-a_{1})(c_{0}-c_{1}))|}$
$=\sqrt{2}\sqrt{|\langle\pi_{\tau}(b_{0})\Omega,\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\rangle|+|\langle\pi_{\tau}(b_{1})\Omega,\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\rangle|}$
$\leq\sqrt{2}\sqrt{\sqrt{\|\pi_{\tau}(b_{0})\Omega\|\|\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\|}+\sqrt{\|\pi_{\tau}(b_{1})\Omega\|\|\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\|}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau((a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2})}+\sqrt{\tau((a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2})}}$
$\leq\sqrt{2}\sqrt{\sqrt{\tau(a_{0}+a_{1})^{2}+\tau(a_{0}-a_{1})^{2}}\sqrt{\tau(c_{0}+c_{1})^{2}+\tau(c_{0}-c_{1})^{2}}}$
$=\sqrt{2}\sqrt{\sqrt{2\tau(a_{0}^{2}+a_{1}^{2})}\sqrt{2\tau(c_{0}^{2}+c_{1}^{2})}}$
$\leq\sqrt{2}\sqrt{2\sqrt{2}\sqrt{2}}$
$=2\sqrt{2}.$
[…]
Below, we demonstrate how the value (3.1) depends on the abelianness of the algebras $\mathcal{M}_{A}$, $\mathcal{M}_{B}$ and $\mathcal{M}_{C}$. The results indicate that in entanglement-swapping networks, the abelianness of the three algebras plays distinct roles in setting the upper bound of the inequality to 2. This is not a simple generalization of the bipartite Bell scenario [17], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. (1) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
(2) If $\mathcal{M}_{B}$ is Abelian, then
$|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, respectively, the eight elements
$a_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),\ \ \ c_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}c_{0})(1+\epsilon_{1}c_{1})$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. By direct computation, one obtains that
$a_{0}+a_{1}=2(a_{++}-a_{--}),\ c_{0}+c_{1}=2(c_{++}-c_{--}),$
$a_{0}-a_{1}=2(a_{+-}-a_{-+}),\ c_{0}-c_{1}=2(c_{+-}-c_{-+}).$
So one obtains that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$=2\left(\sqrt{|\tau((a_{++}-a_{--})b_{0}(c_{++}-c_{--}))|}+\sqrt{|\tau((a_{+-}-a_{-+})b_{1}(c_{+-}-c_{-+}))|}\right)$
$=2\left(\sqrt{|\tau(a_{++}b_{0}c_{++})-\tau(a_{++}b_{0}c_{--})-\tau(a_{--}b_{0}c_{++})+\tau(a_{--}b_{0}c_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(a_{+-}b_{1}c_{+-})-\tau(a_{+-}b_{1}c_{-+})-\tau(a_{-+}b_{1}c_{+-})+\tau(a_{-+}b_{1}c_{-+})|}\right)$
$\leq 2\left(\sqrt{|\tau(a_{++}c_{++})+\tau(a_{++}c_{--})+\tau(a_{--}c_{++})+\tau(a_{--}c_{--})|}\right$.
$\qquad\left.+\sqrt{|\tau(a_{+-}c_{+-})+\tau(a_{+-}c_{-+})+\tau(a_{-+}c_{+-})+\tau(a_{-+}c_{-+})|}\right)$
$=2\left(\sqrt{|\tau(a_{++}+a_{--})||\tau(c_{++}+c_{--})|}+\sqrt{|\tau(a_{+-}+a_{-+})||\tau(c_{+-}+c_{-+})|}\right)$
$\leq 2\sqrt{\tau(a_{++}+a_{--})+\tau(a_{+-}+a_{-+})}\sqrt{\tau(c_{++}+c_{--})+\tau(c_{+-}c_{-+})}$
$=2\sqrt{1}\sqrt{1}=2,$
where the first inequality follows the fact that $I\leq b_{i}\leq I$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $a_{\epsilon_{0}\epsilon_{1}},\ b_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{B}$ is Abelian, the four elements
$b_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. It follows from a direct computation that
$b_{0}+b_{1}=2(b_{++}-b_{--}),\ b_{0}-b_{1}=2(b_{+-}-b_{-+}).$
So one obtains that
$\tau[a_{0}(b_{0}+b_{1})c_{0}]+\tau[a_{1}(b_{0}-b_{1})c_{1}]$
$=$ $2(\tau[a_{0}(b_{++}-b_{--})c_{0}]+\tau[a_{1}(b_{+-}-b_{-+})c_{1}])$
$\leq$ $2(\tau[b_{++}]+\tau[b_{--}]+\tau[b_{+-}]+\tau[b_{-+}])=2$
[…]
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-type inequalities $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ and $|K_{\tau}|+|L_{\tau}|\leq 2$ can indicate the abelianness of the underlying algebras. Conversely, applying the theorem, we can deduce from the violation of an inequality with an upper bound of 2 that the three von Neumann algebras are non-abelian. Here, one also needs to consider the chosen observables and the specific state.
If we wish to set aside the influence of the algebraic structure and consider separately the impact of the choice of observables and the state on values of (3.1) and (3.2), then the following conclusion can be regarded as a generalization of the results in the non-classical mechanical framework. For the sake of clarity, we introduce a tensor product structure into the three-algebra MCvNA model. Let $\mathcal{M}_{B}=\mathcal{M}_{B}^{L}\otimes\mathcal{M}_{B}^{R}$. We call the state $\tau\in\mathcal{M}_{ABC}^{*}$ is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ if there are $\xi_{\alpha}\in\mathcal{M}_{A}^{*},\eta_{\alpha}\in(\mathcal{M}_{B}^{L})^{*}$ such that $\tau(ab^{L})=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a)\eta_{\alpha}(b^{L})$ with $\sum_{\alpha}\lambda_{\alpha}=1$.
Theorem 3. $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ for any state $\tau$ that is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ and $(\mathcal{M}_{B}^{R}\vee\mathcal{M}_{C})^{\prime\prime}$.
Proof. It follows from separability assumption of $\tau$ that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
$=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a_{0}+a_{1})\eta_{\alpha}(b_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(b_{0}^{R})N_{\beta}(c_{0}+c_{1})|}$
$\hskip 18.49988pt+\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a_{0}-a_{1})\eta_{\alpha}(b_{1}^{R})||\sum_{\beta}\mu_{\beta}M_{\beta}(b_{1}^{R})N_{\beta}(c_{0}-c_{1})|}$
$\leq\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(a_{0}+a_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(c_{0}+c_{1})|)}+\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(a_{0}-a_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(c_{0}-c_{1})|)}$
$\leq\sqrt{\sum_{\alpha}\lambda_{\alpha}(|\xi_{\alpha}(a_{0}+a_{1})|+|\xi_{\alpha}(a_{0}-a_{1})|)}\sqrt{\sum_{\beta}\mu_{\beta}(|N_{\beta}(c_{0}+c_{1})|+|N_{\beta}(c_{0}-c_{1})|)}$
$=2\sqrt{\sum_{\alpha}\lambda_{\alpha}\max\{\xi_{\alpha}(a_{0}),\xi_{\alpha}(a_{1})\}}\sqrt{\sum_{\beta}\mu_{\beta}\max\{N_{\beta}(c_{0}),N_{\beta}(c_{1})\}}$
$\leq 2.$
[…]
From the above conclusion, we see that whether a bilocal inequality can be violated depends on the choice of observables and the state, and even more so on the structure of the algebra. In the following, we define a quantity
$\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity (3.2).
Corollary 4.
(1) $2\leq\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})\leq 2\sqrt{2}$.
(2) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(3) $\phi\in(\mathcal{M}_{A}\vee\mathcal{M}_{B_{1}}\vee\mathcal{M}_{B_{2}}\vee\mathcal{M}_{C})^{*}$ is convex combination of separable states, then $\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(4) $|\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})-\mathcal{S}(\psi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})|\leq k\sqrt{\|\phi-\psi\|}$, so, in particular, $\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1). Since $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$. And combinig the proof of Theorem 2, we obtain (1).
(2) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 2.
(3) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 3.
To prove (4). According to the representations of $I_{\tau}$ and $J_{\tau}$, and the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and triangle inequality, one can obtain that
$\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
$=\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\phi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)\right$.
$\qquad\left.-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)\right|$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\phi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right$.
$\qquad\left.-\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right|$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}\right|\right$.
$\qquad\left.+\left|\sqrt{|\phi((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}-\sqrt{|\psi((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right|\right)$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|(\phi-\psi)((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|(\phi-\psi)((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}\right)$
$\leq\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{\|\phi-\psi\|\,\|(a_{0}+a_{1})b_{0}(c_{0}+c_{1})\|}+\sqrt{\|\phi-\psi\|\,\|(a_{0}-a_{1})b_{1}(c_{0}-c_{1})\|}\right)$
$\leq k\sqrt{\|\phi-\psi\|},$
[…]
4. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the inequality, and maximal violation refers to attaining $2\sqrt{2}$.
In the following theorem, we attempt to analyze the conditions for maximal violation.
Theorem 5. If $\tau$ is faithful, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
if and only if $\tau(a_{i}^{2}a)=\tau(a),\ \tau(b_{i}^{2}b)=\tau(b),\ \tau(c_{i}^{2}c)=\tau(c)$, and $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=0,\ \tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},b\in\mathcal{M}_{B},c\in\mathcal{M}_{C}$.
Remark. This theorem states that when the inequality is maximally violated, $I,\ a_{0},\ a_{1}$ and $\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{A}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{A}$. And the same holds for $I,\ c_{0},\ c_{1}$ and $\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{M}_{C}$. Moreover, $\tau(a_{0}a_{1}+a_{1}a_{0})=0,\ \tau(c_{0}c_{1}+c_{1}c_{0})=0.$
Proof. If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 1 that for any $t\in[0,1]$, these equalities hold:
(4.1) $|\tau(b_{0}(a_{0}+a_{1})(c_{0}+c_{1}))|=|\tau(b_{1}(a_{0}-a_{1})(c_{0}-c_{1}))|,$
(4.2) $\pi_{\tau}(b_{0})\Omega=k_{0}\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega,$
(4.3) $\pi_{\tau}(b_{1})\Omega=k_{1}\pi_{\tau}[(a_{0}-a_{1})(c_{0}-c_{1})]\Omega,$
(4.4) $\|\pi_{\tau}(b_{0})\Omega\|=\|\pi_{\tau}(b_{1})\Omega\|=1,$
(4.5) $\tau[(a_{0}+a_{1})^{2}]=t\tau[(c_{0}+c_{1})^{2}],$
(4.6) $\tau[(a_{0}-a_{1})^{2}]=t\tau[(c_{0}-c_{1})^{2}],$
(4.7) $a_{0}^{2}+a_{1}^{2}=2I,c_{0}^{2}+c_{1}^{2}=2I.$
From the (4.7), one gets $a_{i}^{2}=I,\ c_{i}^{2}=I\ (i=0,1)$ because $I\leq a_{i},c_{i}\leq I$, therefore $\tau(a_{i}^{2}a)=\tau(a)\ \tau(c_{i}^{2}c)=\tau(c)$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$.
Now let us show the proof for $\tau(b_{i}^{2}b)=\tau(b),\ \tau[(a_{0}a_{1}+a_{1}a_{0})a]=0$, and $\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$, which implies $b_{i}^{2}=I,\ a_{0}a_{1}+a_{1}a_{0}=0,$ and $c_{0}c_{1}+c_{1}c_{0}=0$. According to the Eqs. (4.5)-(4.7), i.e.,
$\left\{\begin{array}[]{c}\tau(a_{0}^{2}+a_{1}^{2}+a_{0}a_{1}+a_{1}a_{0})=t\tau(c_{0}^{2}+c_{1}^{2}+c_{0}c_{1}+c_{1}c_{0})\\ \tau(a_{0}^{2}+a_{1}^{2}-a_{0}a_{1}-a_{1}a_{0})=t\tau(c_{0}^{2}+c_{1}^{2}-c_{0}c_{1}-c_{1}c_{0})\\ a_{0}^{2}+a_{1}^{2}=2I,\ c_{0}^{2}+c_{1}^{2}=2I,\end{array}\right$.
we can get
$t=1,\ \tau(a_{0}a_{1}+a_{1}a_{0})=\tau(c_{0}c_{1}+c_{1}c_{0}).$
Then combing the conditions (4.2), (4.3) and condition (4.4) one can get
$k_{0}^{2}\tau[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]=1,\ k_{1}^{2}\tau[(a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2}]=1,$
i.e.,
(4.8) $|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$
because of $\tau(X)=\tau(Y)$, where $X=a_{0}a_{1}+a_{1}a_{0},\ Y=c_{0}c_{1}+c_{1}c_{0}$. Then according to Eqs. (4.2), (4.3), and (4.1),
$|\tau[b_{0}(a_{0}+a_{1})(c_{0}+c_{1})]|=|k_{0}||\tau[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]|,$
$|\tau[b_{1}(a_{0}-a_{1})(c_{0}-c_{1})]|=|k_{1}|\tau[(a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2}]|.$
Substituting (4.8) and (4.1) to the above equations, we have
$\frac{1}{2+\tau(X)}(2+\tau(X))^{2}=\frac{1}{2-\tau(X)}(2-\tau(X))^{2}$
deriving that $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(b_{i})\Omega\|=1$ that
$k_{0}^{2}=k_{1}^{2}=\frac{1}{4}$
from the Eq. (4.8). Since
$|\langle\Omega,\pi_{\tau}(b_{0}^{2})\Omega\rangle|$ $=$ $|\frac{1}{4}\langle\Omega,\pi_{\tau}[(a_{0}+a_{1})^{2}(c_{0}+c_{1})^{2}]\Omega\rangle|$
$\leq$ $\frac{1}{4}\sqrt{\|\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega\|}\sqrt{\|\pi_{\tau}[(a_{0}+a_{1})(c_{0}+c_{1})]\Omega\|}$
$=$ $\frac{1}{4}\tau[(a_{0}+a_{1})^{2}]\tau[(c_{0}+c_{1})^{2}]=1,$
and $\tau(b_{0}^{2})=|\langle\Omega,\pi_{\tau}(b_{0}^{2})\Omega\rangle|=1$, this implies that $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$, so
$\tau(b_{0}^{2}b)=\langle\Omega,\pi_{\tau}(b_{0}^{2})\pi_{\tau}(b)\Omega\rangle=\langle\pi_{\tau}(b_{0}^{2})\Omega,\pi_{\tau}(b)\Omega\rangle=\langle\Omega,\pi_{\tau}(b)\Omega\rangle=\tau(b).$
Similarly $\tau(b_{1}^{2}b)=\tau(b)$ for any $b\in\mathcal{M}_{B}.$ Furthermore, it follows from Eq. (4.2) and $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$ that
$\pi_{\tau}(2X+2Y+XY)\Omega=0.$
So $\tau[(2X+2Y+XY)^{2}]=0,$ implying that
$\tau(X^{2})=\tau(Y^{2})=0.$
Combining the faithfulness, non-negativity of the state $\tau$ and the self-adjointness of X,YX,\ Y. Then $X=Y=0$, i.e.,
$a_{0}a_{1}+a_{1}a_{0}=c_{0}c_{1}+c_{1}c_{0}=0,$
and implies that $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ .
[…]
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. However, this does not lead to further global structural properties of the algebra. Below, we aim to further explore the relationship between the maximal violation of the inequality and the global properties of the algebra.
Theorem 6. Let $\mathcal{M}_{ABC}$ be a von Neumann algebra on a separable Hilbert space $H$ with cyclic and separating vector. If there are hyperfinite type II1 factors $\mathcal{R}_{A}$ and $\mathcal{R}_{C}$ such that $\mathcal{M}_{A}\simeq\mathcal{M}_{A}\otimes\mathcal{R}_{A}$ and $\mathcal{M}_{C}\simeq\mathcal{M}_{C}\otimes\mathcal{R}_{C}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
Proof. We first deal with the special case: $\mathcal{M}_{A}=\mathcal{R}_{A}$ and $\mathcal{M}_{C}=\mathcal{R}_{C}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [76], $\mathcal{R}_{A}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now from Theorem 5, we know that $I,\ a_{0},\ a_{1}$ and $a_{2}=-\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $I,\ c_{0},\ c_{1}$ and $c_{2}=-\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{K}_{\alpha}$, and $b_{0},b_{1}$ with $b_{i}^{2}=I$ such that the quantity (3.1) arrives at $2\sqrt{2}$. Applying the similar technologies of Eqs. (6)-(8) of Ref.[72], one can obtain that
(4.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}+a_{1})b_{0}(c_{0}+c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
(4.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}-a_{1})b_{1}(c_{0}-c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}$
[…]
$\widetilde{T}_{\alpha}=(\widetilde{a}_{0,\alpha}+\widetilde{a}_{1,\alpha})\widetilde{b}_{0,\alpha}(\widetilde{c}_{0,\alpha}+\widetilde{c}_{1,\alpha}),$
$\widetilde{S}_{\alpha}=(\widetilde{a}_{0,\alpha}-\widetilde{a}_{1,\alpha})\widetilde{b}_{1,\alpha}(\widetilde{c}_{0,\alpha}-\widetilde{c}_{1,\alpha}).$
and furthermore $\Phi=\bigotimes_{\alpha\in\mathbb{N}}\Phi_{\alpha},\Psi=\bigotimes_{\alpha\in\mathbb{N}}\Psi_{\alpha},$ with $\Phi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Psi_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Phi,\widetilde{T}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2},\sqrt{\langle\Phi,\widetilde{S}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Phi$ and $\Psi$).
[…]
It is mentioned that the results in Theorem 6 can be applied into the quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions are typically type III1 factors (a classification indicating the algebra has no trace and no pure normal states, reflecting the infinite degrees of freedom of QFT). By the classical result of A. Connes, all injective (i.e., amenable or hyperfinite, a property related to physical phase space compactness) infinite factors (excepting some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\otimes\mathcal{R}_{1}$). So the assumption of Theorem 6 is satisfied.
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If each $\mathcal{A}(\mathcal{O}_{i})$ is III1 factor, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. All authors contributed equally to this work.

### v2

@@
-arXiv:2603.01466v1[math.FA] 02 Mar 2026
+Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models
 Abstract.
-Differently from the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states.
-Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework often reveals information about the algebraic structure.
+In contrast to the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states.
+Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework can be applied to addressing questions in quantum field theory, while also providing insights into the structural properties of von Neumann algebras.
 In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities.
@@
 Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse.
-Our results not only utilize the violation of bilocal inequalities to reveal the structural properties of von Neumann algebras, but can also be applied to quantum mechanics and quantum field theory.
 Corresponding author
@@
 1. Introduction
-Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3].
+In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3].
 Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8].
 It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
-Moreover, Bell nonlocality in quantum field theory (QFT) serves as a crucial bridge connecting quantum information science with fundamental physics [16, 17].
-It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [19, 18].
-Furthermore, the tension between nonlocal quantum phenomena and the locality assumptions of general relativity offers profound insights for quantum gravity research, suggesting that spacetime itself may be an emergent property rather than a fundamental aspect of reality [20, 21, 22].
+However, it should be pointed out that relying solely on the non-relativistic quantum mechanics to discuss quantum information problems has two drawbacks [16, 17, 18]: 1) it fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom; 2) it is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras.
+Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results [19, 20, 21, 22, 23, 24, 25].
+In the 1980s early, mathematician Summers et al.
+introduce firstly the maximal violation of Bell inequalities and prove that its value is bounded by $2\sqrt${2} in the von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [26].
+This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [27].
+Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [28, 29, 30, 31, 32].
+These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [27].
+This is a pioneering work to make Bell nonlocality in quantum field theory (QFT) serve as a crucial bridge connecting quantum information science with fundamental physics [27, 33].
+It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [34, 35, 36, 37, 38].
+In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states.
+Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [39, 40, 41, 42, 43, 44].
+To date, Bell-like inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [39, 45], chain configurations [46, 47], star topologies [48, 49], polygon structures [50, 51, 52], tree-shaped networks [53, 54, 55], arbitrary acyclic networks [41, 42, 56], and arbitrary $k$-independent networks [57].
+Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [58, 59].
+Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [60].
+Recently, the notion of bi-locality in an entanglement swapping network based on the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al.
+[61], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [62].
+However, Bell-like inequalities for such von Neumann algebraic models have not yet been established.
+In this paper, we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
+2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
 In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [63, 64, 65].
-This differs from the traditional quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [27, 26, 28].
+This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [66, 67, 68].
 These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model.
-In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, denoted as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators.
+In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{B}}$, denoted as $\mathcal{H}=\mathcal{H}_{\mathcal{A}}\otimes\mathcal{H}_{\mathcal{B}}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators.
 Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$.
-A quantum state $\rho$ is a positive operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$.
-The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ However, it should be pointed out that relying solely on the above tensor-product model to discuss quantum information problems has two drawbacks [29, 30, 31]: 1) It fails to provide a universal framework for accurately describing phenomena in infinite-dimensional systems.
-2) It is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras.
-In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, where $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$, and $\mathcal{M}_{A},\mathcal{M}_{B}$ are two mutually commuting von Neumann subalgebras of $\mathcal{M}$.
-The notation $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
+A quantum state $\rho$ is a positive semidefinite operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$.
+The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$.
+Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
 However, in more general cases, a natural question arises: are the two models isomorphic?
-In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{A}\otimes\mathcal{M}_{B}$?
+In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}$?
 This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [69, 70, 71].
@@
 In 2020, Ji and collaborators [71] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative.
-This means that the tensor-product model algebra model and the mutually-commuting von Neumann algebra model are not equivalent.
+This means that the tensor-product algebra model and the mutually-commuting von Neumann algebra model are not equivalent.
 Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [71, 72, 73, 74].
-Overall, Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [17].
-Summers et al.
-introduces the maximal Bell correlation and proves that its value is bounded by $\sqrt{2}$ in mutually-commuting von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [38].
-Translating these bounds into the vacuum representation of algebraic quantum field theory, he shows that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges and decay exponentially with mass-governed distance [39, 40, 41, 42, 43].
-These works expose a new algebraic invariant¡ªthe Bell correlation invariant¡ªthat distinguishes infinitely many isomorphism classes of mutually-commuting von Neumann algebra pairs and links maximal violation to the presence of the hyperfinite type-${\rm II}_{1}$ factor [17].
-In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states.
-Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [44, 45, 46, 47, 48, 49].
-To date, Bell-type inequalities have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [44, 50], chain configurations [51, 52], star topologies [53, 54], polygon structures [55, 56, 57], tree-shaped networks [58, 59, 60], arbitrary acyclic networks [46, 47, 61], and arbitrary $k$-independent networks [62].
-Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [63, 64].
-Furthermore, the interplay between graph theory and quantum correlations has attracted considerable attention [65, 66].
-Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum topologies [67].
-Interestingly, the notion of bi-locality in the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al [68], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [69].
-In this paper we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
 3. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
@@
 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them.
-All the three parties can perform measurements (or inputs) on their system, labeled $x$, $y$ and $z$ for Alice, Bob and Charles.
-After measuring the three parties, they obtain outcomes denoted $a$, $b$ and $c$, respectively.
-Here, $x,y,z,a,b,c\in\{0,1\}$.
-Bob’s measurement $y$ might correspond to a joint measurement on the two systems that he receives from each source.
+The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively, where $x,y,z,a,b,c\in\{0,1\}$.
+Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively.
+Here it is required that the spectra of operators $A_{x},B_{y},C_{Z}$ are all {-1,1}, implying that $I\leq A_{x},B_{y},C_{z}\leq I$.
 The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$.
-In this scenario, the joint probability distribution $p(abc|xyz)$ is bilocal if it can be written as
-$p(abc|xyz)=\int\int d\lambda d\mu q_{1}(\lambda)q_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
-where $\lambda$ ($\mu$) characterizes the joint hidden variable of the system produced by the source $\rho_{AB}$ ($\rho_{BC}$) [70, 71].
-Let $A_{x},B_{y},C_{z}$ denote the measurements of Alice, Bob, and Charles with inputs $x$, $y$, $z$, respectively.
-Note that the middle party Bob always performs joint measurement.
-Then, the bilocal inequality can be written as:
+In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
+$p(abc|xyz)=\int\int d\lambda d\mu p_{1}(\lambda)p_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
+where $\lambda$ and $\mu$ characterize the hidden variables of the systems produced by the sources $\rho_{AB}$ and $\rho_{BC}$, respectively [75, 76]. Otherwise, it is called non-bilocal.
+In order to detect non-bilocal correlations generated by the network, it is often necessary to find suitable measurements that violate the following bilocal inequality
 (3.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
-where
+whose maximum quantum violation is $2\sqrt{2}$ and is attainable. Here
 $I\equiv\sum_{x,z}\langle A_{x}B_{0}C_{z}\rangle=\langle(A_{0}+A_{1})B_{0}(C_{0}+C_{1})\rangle,$
 $J\equiv\sum_{x,z}(-1)^{x+z}\langle A_{x}B_{1}C_{z}\rangle=\langle(A_{0}-A_{1})B_{1}(C_{0}-C_{1})\rangle$
-and $\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz)$ [72].
-In addition, one can choose different observables to construct different bilocal inequalities. For instance, the violation of the inequality
+as introduced in Ref. [77]:
+$\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
+Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurement (POVMs) performed by Alice, Bob and Charles, respectively.
+Furthermore, it follows from Ref. [57] that the violation of the inequality
 (3.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
 can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [78].
-Of course, different inequalities exhibit varying capabilities in detecting nonlocality.
-Mutually-commuting von Neumann algbera models.
-In QFT, the observables for Alice, Bob and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{A},\ \mathcal{M}_{B},\ \mathcal{M}_{C}$.
-In general, they are all required to be type III von Neumann algebra factors.
-Recall two von Neumann algebras satisfy Haag duality means that each algebra equals the commutant of the other.
-Haag duality encode quantum nonlocality intrinsically within their infinite-dimensional non-abelian structure through their type III factor property, which enforces that local algebras of spacelike separated regions are entangled with trivial relative commutants [74, 75].
-Here, we do not restrict von Neumann algebras to III type factors, and generalize the Haag duality condition to mutual commutativity, proposing a von Neumann algebra model with mutual commutativity.
-Therefore, our model covers both scenarios of non-relativistic quantum mechanics and quantum field theory.
+Notably, different inequalities exhibit varying capabilities in detecting nonlocality.
+Mutually-commuting von Neumann algebra models.
+In QFT, the observables for Alice, Bob, and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, $\mathcal{M}_{\mathcal{C}}$.
+Therefore, our model encompasses both the non-relativistic quantum mechanics scenario and the quantum field theory scenario.
 The idea of this model is similar to that in Refs.
@@
 Definition 1.
-(Ternary mutually-commuting von Neumann algebra models) Let $\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C}$ be von Neumann subalgebras of $\mathcal{B}(H)$ over some Hilbert space $H$, and they are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{A,B,C\}$, where $\mathcal{M}_{j}^{\prime}$ is the communicator of $\mathcal{M}_{j}$.
+(Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$.
 The generated von Neumann algebra
-$\mathcal{M_{ABC}}=(\mathcal{M_{A}}\vee\mathcal{M_{B}}\vee\mathcal{M_{C}})^{\prime\prime}.$
-We refer to the above model as the MCvNA model.
-When $\mathcal{M}_{ABC}=\mathcal{M}_{A}\otimes\mathcal{M}_{B}\otimes\mathcal{M}_{C}$, it is called the tensor product algebra model.
-Here, in order to correspond to the Hermitian property of quantum observables, these three algebras are always restricted to consist of self-adjoint elements.
-We point out that the von Neumann algebra model defined above is intended for the study of algebraic structural properties; therefore, we have adopted the most general definition, which encompasses but does not correspond to models of quantum mechanics or quantum field theory.
-We intend to use the above model of three mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig.
+$\mathcal{M}_{\mathcal{ABC}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}.$
+We refer to the above model as the TMCvNA model.
+When $\mathcal{M}_{\mathcal{ABC}}\simeq\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, it is called the tensor product algebra model.
+In this paper, for any $A\in\mathcal{M}_{\mathcal{A}}$, $B\in\mathcal{M}_{\mathcal{B}}$, $C\in\mathcal{M}_{\mathcal{C}}$, we always assume that they are Hermitian.
+We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig.
 1.
-We note that there is no correlation between parties $A$ and $C$ in the network.
+We note that there is no correlation between the parties $\mathcal{A}$ and $\mathcal{C}$ in the network.
 Mathematically, this independence can be described by the following formula.
-Let $\tau$ be a state in the dual space $\mathcal{M}_{ABC}^{*}$.
-In correspondence with the bilocal scenario, we always require that
+So the network state of the entanglement swapping network $\tau$ should be a state in the dual space $\mathcal{M}_{\mathcal{ABC}}^{*}$, satisfying
 () $\tau(AC)=\tau(A)\tau(C)$
-for any $a\in\mathcal{M}_{A},c\in\mathcal{M}_{C}$. We call it as the independent condition (*). This assumption will be used throughout this paper.
+for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call it the independent condition (*). This assumption will be used throughout this paper.
+Definition 2.
+A ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with the independence condition (*).
 4. Bilocal inequalities and their bounds
-Following the assumptions in Section 2, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ is defined as
+In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network.
+Specifically, in this model, we can construct bilocal inequalities analogous to those in the non-relativistic setting.
+Define
+$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),\ J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
+where $\tau$ is the state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*).
+Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively.
+Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as
 $p(\alpha\beta\gamma|xyz)=\tau({A}_{\alpha|x}{B}_{\beta|y}{C}_{\gamma|z}),$
-where the measurement operators ${A}_{\alpha|x}\in\mathcal{M}_{A},\ {B}_{\beta|y}\in\mathcal{M}_{B},\ {C}_{\gamma|z}\in\mathcal{M}_{C}$, $\{x,y,z\}$ and $\{\alpha,\beta,\gamma\}$ denote the inputs and outputs of measurement operators with $x,y,z,\alpha,\beta,\gamma\in\{0,1\}$.
-Let $a_{x}\in\mathcal{M}_{A}$, $b_{y}\in\mathcal{M}_{B}$, $c_{z}\in\mathcal{M}_{C}$ be operators such that $I\leq a_{x},b_{y},c_{z}\leq I$ for any $x,y,z\in\{0,1\}$, and $a_{x}=\sum_{\alpha=0}^{1}(-1)^{\alpha}A_{\alpha|x},b_{y}=\sum_{\beta=0}^{1}(-1)^{\beta}B_{\beta|y},c_{z}=\sum_{\gamma=0}^{1}(-1)^{\gamma}C_{\gamma|z}$.
-Then denote $I_{\tau}=\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1})),\ J_{\tau}=\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))$.
-In the MCvNA model, analogous to Ineq.
-(2.1), we define
+In the TMCvNA model, analogous to Ineq. (3.1), we set
 (4.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
-Analogous to Ineq. (2.2), under the same assumption, we have the following
-(3.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|.$
+Analogously, corresponding to Ineq. (3.2), we define
+(4.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|,$
 where, $K_{\tau}=\tau(A_{0}(B_{0}+B_{1})C_{0}),\ L_{\tau}=\tau(A_{1}(B_{0}-B_{1})C_{1})$.
+We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$.
+The same criterion applies to $\mathcal{S}_{\tau}^{\prime}$ in Ineq.
+(4.2).
 Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively.
-We mention that all theorems in this section are based on the independence assumption (*).
-The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the maximum upper bound that the quantity (3.1) can achieve is $2\sqrt{2}$.
-This is analogous to the case in non-relativistic quantum mechanics, where the maximal violation of the bilocal value based on quantum resources is $2\sqrt{2}$.
-Theorem 1. In the MCvNA model, $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
+The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the supremum of $\mathcal{S}_{\tau}$ defined in Ineq.
+(4.1) is $2\sqrt{2}$.
+This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq.
+(3.1) attains a maximal violation of $2\sqrt{2}$ based on quantum resources.
+Theorem 1. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
 Proof.
-According to Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{ABC}\rightarrow H_{\tau}$ and a cyclic vector $\Omega\in H_{\tau}$ such that the set $\{\pi_{\tau}(d)\Omega:\ d\in\mathcal{M}_{ABC}\}$ is dense in $H_{\tau}$.
+According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{B}(\mathcal{H}_{\tau})$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$.
 It follows by applying the Cauchy-Schwarz inequality that
-$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
 $=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
@@
 $=\sqrt{2}\sqrt{|\langle\pi_{\tau}(B_{0})\Omega,\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\rangle|+|\langle\pi_{\tau}(B_{1})\Omega,\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\rangle|}$
-$\leq\sqrt{2}\sqrt{\sqrt{\|\pi_{\tau}(b_{0})\Omega\|\|\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\|}+\sqrt{\|\pi_{\tau}(b_{1})\Omega\|\|\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\|}}$
+$\leq\sqrt{2}\sqrt{\|\pi_{\tau}(B_{0})\Omega\|\|\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\|+\|\pi_{\tau}(B_{1})\Omega\|\|\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\|}$
 $\leq\sqrt{2}\sqrt{\sqrt{\tau((A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2})}+\sqrt{\tau((A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2})}}$
@@
 […]
-Below, we demonstrate how the value (3.1) depends on the abelianness of the algebras $\mathcal{M}_{A}$, $\mathcal{M}_{B}$ and $\mathcal{M}_{C}$.
-The results indicate that in entanglement-swapping networks, the abelianness of the three algebras plays distinct roles in setting the upper bound of the inequality to 2.
+Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Ineq.
+(4.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$.
+Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur.
 This is not a simple generalization of the bipartite Bell scenario [27], where, with only two systems, Summers et al.
 showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
-Theorem 2. (1) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then
-$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
-(2) If $\mathcal{M}_{B}$ is Abelian, then
-$|K_{\tau}|+|L_{\tau}|\leq 2.$
-Proof. To see (1), since $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, respectively, the eight elements
+Theorem 2. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
+(1) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2;$
+(2) if $\mathcal{M}_{\mathcal{B}}$ is Abelian, then
+$\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|\leq 2.$
+Proof. To see (1), since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
 $A_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}A_{0})(1+\epsilon_{1}A_{1}),\ \ \ C_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}C_{0})(1+\epsilon_{1}C_{1})$
@@
 So one obtains that
-$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
 $=2\left(\sqrt{|\tau((A_{++}-A_{--})B_{0}(C_{++}-C_{--}))|}+\sqrt{|\tau((A_{+-}-A_{-+})B_{1}(C_{+-}-C_{-+}))|}\right)$
@@
 $=2\sqrt{1}\sqrt{1}=2,$
-where the first inequality follows the fact that $I\leq b_{i}\leq I$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $a_{\epsilon_{0}\epsilon_{1}},\ b_{\epsilon_{0}\epsilon_{1}}$.
-To prove (2), since $\mathcal{M}_{B}$ is Abelian, the four elements
+where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ B_{\epsilon_{0}\epsilon_{1}}$.
+To prove (2), since $\mathcal{M}_{\mathcal{B}}$ is Abelian, the four elements
 $B_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
@@
 So one obtains that
-$\tau[a_{0}(b_{0}+b_{1})c_{0}]+\tau[a_{1}(b_{0}-b_{1})c_{1}]$
+$\mathcal{S}_{\tau}^{\prime}$ $=$ $\tau[A_{0}(B_{0}+B_{1})C_{0}]+\tau[A_{1}(B_{0}-B_{1})C_{1}]$
 $=$ $2(\tau[A_{0}(B_{++}-B_{--})C_{0}]+\tau[A_{1}(B_{+-}-B_{-+})C_{1}])$
@@
 Remark.
-The above theorem illustrates a phenomenon: the violations of the two Bell-type inequalities $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ and $|K_{\tau}|+|L_{\tau}|\leq 2$ can indicate the abelianness of the underlying algebras.
-Conversely, applying the theorem, we can deduce from the violation of an inequality with an upper bound of 2 that the three von Neumann algebras are non-abelian.
+The above theorem illustrates a phenomenon: the violations of the two Bell-like inequalities, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ and $2<\mathcal{S}^{\prime}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras.
+Applying the theorem, we can infer from a violation of Ineq.
+(4.1) that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian; Similarly, a violation of Ineq.
+(4.2) implies that $\mathcal{M_{B}}$ is non-abelian.
 Here, one also needs to consider the chosen observables and the specific state.
-If we wish to set aside the influence of the algebraic structure and consider separately the impact of the choice of observables and the state on values of (3.1) and (3.2), then the following conclusion can be regarded as a generalization of the results in the non-classical mechanical framework.
-For the sake of clarity, we introduce a tensor product structure into the three-algebra MCvNA model.
-Let $\mathcal{M}_{B}=\mathcal{M}_{B}^{L}\otimes\mathcal{M}_{B}^{R}$.
-We call the state $\tau\in\mathcal{M}_{ABC}^{*}$ is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ if there are $\xi_{\alpha}\in\mathcal{M}_{A}^{*},\eta_{\alpha}\in(\mathcal{M}_{B}^{L})^{*}$ such that $\tau(ab^{L})=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a)\eta_{\alpha}(b^{L})$ with $\sum_{\alpha}\lambda_{\alpha}=1$.
-Theorem 3.
-$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ for any state $\tau$ that is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ and $(\mathcal{M}_{B}^{R}\vee\mathcal{M}_{C})^{\prime\prime}$.
-Proof. It follows from separability assumption of $\tau$ that
-$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
+Nevertheless, it is important to note that these conclusions depend not only on the algebraic structure but also on the specific choice of observables and the particular state under consideration.
+To isolate the influence of the algebraic structure and examine separately the role of observables and the state in determining the values $\mathcal{S}_{\tau}$ and $\mathcal{S}_{\tau}^{\prime}$ in Ineqs.
+(4.1) and  (4.2), the following conclusion can be regarded as a generalization of the results in the non-relativistic quantum mechanical framework.
+For the sake of clarity, we introduce a tensor product structure into the TMCvNA model.
+Before presenting the following result, we first introduce the definition of separable states.
+Definition 3.
+(Separable states) Let $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ be mutually commuting von Neumann subalgebras of $\mathcal{M}_{\mathcal{AB}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}$.
+A state $\tau\in\mathcal{M}_{\mathcal{AB}}^{*}$ is said to be separable if there exist states $\xi_{\alpha}\in\mathcal{M}_{\mathcal{A}}^{*}$, $\eta_{\alpha}\in\mathcal{M}_{\mathcal{B}}^{*}$, and positive coefficients $\lambda_{\alpha}$ with $\sum_{\alpha}\lambda_{\alpha}=1$ such that for all $A\in\mathcal{M}_{\mathcal{A}},B\in\mathcal{M}_{\mathcal{B}}$,
+$\tau(AB)=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A)\eta_{\alpha}(B).$
+[…]
+$\sup_{\{A_{x},B_{y},C_{z}\}}\mathcal{S}_{\tau}=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})=2.$
+[…]
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
 $=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}+A_{1})\eta_{\alpha}(B_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{0}^{R})N_{\beta}(C_{0}+C_{1})|}$
@@
 In the following, we define a quantity
-$\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
-Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity (3.2).
-Corollary 4.
-(1) $2\leq\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})\leq 2\sqrt{2}$.
-(2) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
-(3) $\phi\in(\mathcal{M}_{A}\vee\mathcal{M}_{B_{1}}\vee\mathcal{M}_{B_{2}}\vee\mathcal{M}_{C})^{*}$ is convex combination of separable states, then $\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
-(4) $|\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})-\mathcal{S}(\psi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})|\leq k\sqrt{\|\phi-\psi\|}$, so, in particular, $\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
-Proof. To show (1). Since $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$. And combinig the proof of Theorem 2, we obtain (1).
-(2) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 2.
-(3) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 3.
-To prove (4). According to the representations of $I_{\tau}$ and $J_{\tau}$, and the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and triangle inequality, one can obtain that
-$\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
+$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})$
+that appears in Theorem 3.
+Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity $\mathcal{S}^{\prime}_{\tau}$ in Ineq. (4.2).
+Corollary 4. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
+(1) for any state $\tau$ and any choice of observables $A_{x}\in\mathcal{M}_{\mathcal{A}},\ B_{y}\in\mathcal{M}_{\mathcal{B}},\ C_{z}\in\mathcal{M}_{\mathcal{C}}$ in a scheme with two inputs and two outputs, the quantity $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})$ satisfies
+$2\leq\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})\leq 2\sqrt{2}.$
+(2) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=2$.
+(3) for any states $\phi,\psi\in[(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}]^{*}$, the following inequality holds:
+$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|\leq k\sqrt{\|\phi-\psi\|},$
+where $k$ is a positive constant.
+Consequently, the functional $\phi\rightarrow\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
+Proof. To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 2, we obtain (1).
+(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 2.
+To prove (3), note that by the representations of $I_{\tau}$ and $J_{\tau}$, together with the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and the triangle inequality, one can obtain
+$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|$
+$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
 $=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|\phi((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\phi((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}\right)\right$.
@@
 In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra.
-Here, violation refers to exceeding the upper bound of 2 for the inequality, and maximal violation refers to attaining $2\sqrt{2}$.
+Here, violation refers to exceeding the upper bound of 2 for the $\mathcal{S}_{\tau}$ in inequality (4.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
 In the following theorem, we attempt to analyze the conditions for maximal violation.
-Theorem 5. If $\tau$ is faithful, then
+Theorem 5. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\tau$ is faithful, then
 $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
-if and only if $\tau(a_{i}^{2}a)=\tau(a),\ \tau(b_{i}^{2}b)=\tau(b),\ \tau(c_{i}^{2}c)=\tau(c)$, and $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=0,\ \tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},b\in\mathcal{M}_{B},c\in\mathcal{M}_{C}$.
+if and only if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ with $i\in\{0,1\}$.
 Remark.
-This theorem states that when the inequality is maximally violated, $I,\ a_{0},\ a_{1}$ and $\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{A}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{A}$.
-And the same holds for $I,\ c_{0},\ c_{1}$ and $\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{M}_{C}$.
+This theorem states that when the inequality is maximally violated, $A_{0},\ A_{1}$ and $\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{\mathcal{A}}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{\mathcal{A}}$.
+The same holds for $C_{0},\ C_{1}$ and $\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{M}_{\mathcal{C}}$.
 Moreover, $\tau(A_{0}A_{1}+A_{1}A_{0})=0,\ \tau(C_{0}C_{1}+C_{1}C_{0})=0.$
@@
 (5.7) $A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$
-From the (4.7), one gets $a_{i}^{2}=I,\ c_{i}^{2}=I\ (i=0,1)$ because $I\leq a_{i},c_{i}\leq I$, therefore $\tau(a_{i}^{2}a)=\tau(a)\ \tau(c_{i}^{2}c)=\tau(c)$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$.
-Now let us show the proof for $\tau(b_{i}^{2}b)=\tau(b),\ \tau[(a_{0}a_{1}+a_{1}a_{0})a]=0$, and $\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$, which implies $b_{i}^{2}=I,\ a_{0}a_{1}+a_{1}a_{0}=0,$ and $c_{0}c_{1}+c_{1}c_{0}=0$.
-According to the Eqs.
+From (5.7), one gets $A_{i}^{2}=I,\ C_{i}^{2}=I\ (i=0,1)$ because $I\leq A_{i},C_{i}\leq I$, therefore $\tau(A_{i}^{2}A)=\tau(A),\ \tau(C_{i}^{2}C)=\tau(C)$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$.
+Now, let us show the proof for $\tau(B_{i}^{2}B)=\tau(B),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, and $\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$, which implies $B_{i}^{2}=I,$ $A_{0}A_{1}+A_{1}A_{0}=0,$ and $C_{0}C_{1}+C_{1}C_{0}=0$.
+According to Eqs.
 (5.5)-(5.7), i.e.,
@@
 $t=1,\ \tau(A_{0}A_{1}+A_{1}A_{0})=\tau(C_{0}C_{1}+C_{1}C_{0}).$
-Then combing the conditions (4.2), (4.3) and condition (4.4) one can get
+Then combing conditions (5.2), (5.3) and condition (5.4), one can get
 $k_{0}^{2}\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]=1,\ k_{1}^{2}\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]=1,$
@@
 $|\tau[B_{0}(A_{0}+A_{1})(C_{0}+C_{1})]|=|k_{0}||\tau[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]|,$
-$|\tau[b_{1}(a_{0}-a_{1})(c_{0}-c_{1})]|=|k_{1}|\tau[(a_{0}-a_{1})^{2}(c_{0}-c_{1})^{2}]|.$
+$|\tau[B_{1}(A_{0}-A_{1})(C_{0}-C_{1})]|=|k_{1}||\tau[(A_{0}-A_{1})^{2}(C_{0}-C_{1})^{2}]|.$
 Substituting (5.8) and (5.1) to the above equations, we have
 $\frac{1}{2+\tau(X)}(2+\tau(X))^{2}=\frac{1}{2-\tau(X)}(2-\tau(X))^{2}$
-deriving that $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(b_{i})\Omega\|=1$ that
+deriving $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(B_{i})\Omega\|=1$ that
 $k_{0}^{2}=k_{1}^{2}=\frac{1}{4}$
-from the Eq. (4.8). Since
+from Eq. (5.8). Since
 $|\langle\Omega,\pi_{\tau}(B_{0}^{2})\Omega\rangle|$ $=$ $|\frac{1}{4}\langle\Omega,\pi_{\tau}[(A_{0}+A_{1})^{2}(C_{0}+C_{1})^{2}]\Omega\rangle|$
@@
 $\tau(B_{0}^{2}B)=\langle\Omega,\pi_{\tau}(B_{0}^{2})\pi_{\tau}(B)\Omega\rangle=\langle\pi_{\tau}(B_{0}^{2})\Omega,\pi_{\tau}(B)\Omega\rangle=\langle\Omega,\pi_{\tau}(B)\Omega\rangle=\tau(B).$
-Similarly $\tau(b_{1}^{2}b)=\tau(b)$ for any $b\in\mathcal{M}_{B}.$ Furthermore, it follows from Eq. (4.2) and $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$ that
+Similarly $\tau(B_{1}^{2}B)=\tau(B)$ for any $B\in\mathcal{M}_{\mathcal{B}}.$ Furthermore, it follows from Eq. (5.2) and $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$ that
 $\pi_{\tau}(2X+2Y+XY)\Omega=0.$
@@
 The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor.
-However, this does not lead to further global structural properties of the algebra.
-Below, we aim to further explore the relationship between the maximal violation of the inequality and the global properties of the algebra.
+In contrast, the follwing theorem presents a highly specialized scenario in which a precise characterization of algebras attaining maximal viloation can be obtained.
+Note that next we denote by $\Lambda_{1}\overline{\otimes}\Lambda_{2}$ the von Neumann algebra of the tensor product of $\Lambda_{1},\Lambda_{2}$
 Theorem 6.
-Let $\mathcal{M}_{ABC}$ be a von Neumann algebra on a separable Hilbert space $H$ with cyclic and separating vector.
-If there are hyperfinite type II1 factors $\mathcal{R}_{A}$ and $\mathcal{R}_{C}$ such that $\mathcal{M}_{A}\simeq\mathcal{M}_{A}\otimes\mathcal{R}_{A}$ and $\mathcal{M}_{C}\simeq\mathcal{M}_{C}\otimes\mathcal{R}_{C}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
+In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if there are hyperfinite type II1 factors $\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{C}}$ such that $\mathcal{M}_{\mathcal{A}}\simeq\mathcal{M}_{\mathcal{A}}\overline{\otimes}\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}\simeq\mathcal{M}_{\mathcal{C}}\overline{\otimes}\mathcal{R}_{\mathcal{C}}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
 Proof.
-We first deal with the special case: $\mathcal{M}_{A}=\mathcal{R}_{A}$ and $\mathcal{M}_{C}=\mathcal{R}_{C}$.
-Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [76], $\mathcal{R}_{A}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$.
-Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$.
-We can obtain that there is $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$.
-Now from Theorem 5, we know that $I,\ a_{0},\ a_{1}$ and $a_{2}=-\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $I,\ c_{0},\ c_{1}$ and $c_{2}=-\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{K}_{\alpha}$, and $b_{0},b_{1}$ with $b_{i}^{2}=I$ such that the quantity (3.1) arrives at $2\sqrt{2}$.
-Applying the similar technologies of Eqs.
+We first deal with the special case: $\mathcal{M}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{R}_{\mathcal{C}}$.
+Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [79], $\mathcal{R}_{\mathcal{A}}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$.
+Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$.
+We can obtain that there is $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$.
+Now, from Theorem 5, we know that $A_{0},\ A_{1}$ and $A_{2}=-\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $C_{0},\ C_{1}$ and $C_{2}=-\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{K}_{\alpha}$, and $B_{0},B_{1}$ with $B_{i}^{2}=I$ such that the quantity $\mathcal{S}_{\tau}$ in Ineq.
+(4.1) arrives at $2\sqrt{2}$.
+Applying similar technologies of Eqs.
 (6)-(8) of Ref.[77], one can obtain that
 (5.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}+A_{1})B_{0}(C_{0}+C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
-(4.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}-a_{1})b_{1}(c_{0}-c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}$
+(5.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}-A_{1})B_{1}(C_{0}-C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}.$
 […]
@@
 $\widetilde{S}_{\alpha}=(\widetilde{A}_{0,\alpha}-\widetilde{A}_{1,\alpha})\widetilde{B}_{1,\alpha}(\widetilde{C}_{0,\alpha}-\widetilde{C}_{1,\alpha}).$
-and furthermore $\Phi=\bigotimes_{\alpha\in\mathbb{N}}\Phi_{\alpha},\Psi=\bigotimes_{\alpha\in\mathbb{N}}\Psi_{\alpha},$ with $\Phi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Psi_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$.
-By definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$.
+Furthermore, $\Xi=\bigotimes_{\alpha\in\mathbb{N}}\Xi_{\alpha},\Upsilon=\bigotimes_{\alpha\in\mathbb{N}}\Upsilon_{\alpha},$ with $\Xi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Upsilon_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$.
+By the definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$.
 But one has
-$\sqrt{\langle\Phi,\widetilde{T}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2},\sqrt{\langle\Phi,\widetilde{S}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2}.$
-for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Phi$ and $\Psi$).
+$\sqrt{\langle\Xi,\widetilde{T}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2},\sqrt{\langle\Xi,\widetilde{S}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2}.$
+for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Xi$ and $\Upsilon$).
 […]
 […]
-It is mentioned that the results in Theorem 6 can be applied into the quantum field theory.
-In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions are typically type III1 factors (a classification indicating the algebra has no trace and no pure normal states, reflecting the infinite degrees of freedom of QFT).
+It is mentioned that the results of Theorem 6 can be applied to quantum field theory.
+In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions, are typically type III1 factors.
 By the classical result of A.
-Connes, all injective (i.e., amenable or hyperfinite, a property related to physical phase space compactness) infinite factors (excepting some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\otimes\mathcal{R}_{1}$).
-So the assumption of Theorem 6 is satisfied.
+Connes, all injective infinite factors (except some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\overline{\otimes}\mathcal{R}_{1}$).
 Corollary 7.
-Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector.
-If each $\mathcal{A}(\mathcal{O}_{i})$ is III1 factor, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
-Acknowledgement. All authors contributed equally to this work.
+Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ be the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector.
+If $\mathcal{A}(\mathcal{O}_{1})$ and $\mathcal{A}(\mathcal{O}_{3})$ are both type III1 factors, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
+Acknowledgement. Authors thank for your comments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271394).

### v5

@@
-Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models
-Abstract.
-In contrast to the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states.
-Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework can be applied to addressing questions in quantum field theory, while also providing insights into the structural properties of von Neumann algebras.
-In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities.
-We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras.
-Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse.
-Corresponding author
-Key words and phrases: von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
-1. Introduction
+B. Zheng, S. Yang, J. Hou and K. He
+[…]
+[…]
+[…]
+B. Zheng, S. Yang, J. Hou and K. He
+[…]
+Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models
+College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
+School of Mathematics, North University of China, Taiyuan, 030051, China,
+School of Cyberspace Science and Technology, Beijing Institute of Technology, Beijing, 100081, China
+E-mail :: yangshuyuan2000@163.com
+College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
+E-mail :: jinchuanhou@aliyun.com
+College of Mathematics, Taiyuan University of Technology, Taiyuan, 030024, China
+E-mail :: hekan@tyut.edu.cn
+Violation of Bell-type Inequalities on Mutually-commuting von Neumann Algebra Models of Entanglement Swapping Networks  [^1] Shuyuan Yang and Bingke Zheng contribute equally to the work.
+Supported by the National Natural Science Foundation of China (Grant No.
+12271394, 12571138)
+Bingke ZHENG    Shuyuan YANG    Jinchuan HOU    Kan HE
+[…]
+Violation of Bell inequalities in bipartite systems represented by mutually-commuting von Neumann algebras has pioneered the study of vacuum entanglement in algebraic quantum field theory.
+It is unexpected that the maximal violation of Bell inequality can discover algebraic structures.
+In the paper, we establish the mutually-commuting von Neumann algebra model for entanglement swapping networks and Bell-type inequalities on this model.
+It generalizes the bipartite case to the ternary case.
+These algebras are all general von Neumann algebras, which provide a natural perspective to investigate Bell nonlocality in quantum networks in the infinitely-many-degree-of-freedom setting.
+We determine various bounds for Bell-type inequalities based on the structure of von Neumann algebras, and identify the algebraic structural conditions required for their violation.
+Finally, we show that the maximal violation of Bell-type inequalities in entanglement swapping networks can be used to determine partially the type classification of the underlying von Neumann algebras.
+[…]
+von Neumann algebra, Bell nonlocality, Bell inequality, Quantum network
+[…]
+81R15, 47A63, 47L30, 47N50
+1 Introduction
 In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3].
@@
 It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
-However, it should be pointed out that relying solely on the non-relativistic quantum mechanics to discuss quantum information problems has two drawbacks [16, 17, 18]: 1) it fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom; 2) it is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras.
-Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results [19, 20, 21, 22, 23, 24, 25].
-In the 1980s early, mathematician Summers et al.
-introduce firstly the maximal violation of Bell inequalities and prove that its value is bounded by $2\sqrt${2} in the von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [26].
+Meanwhile, motivated by the quantum field theory (QFT), which originates from the study of relativistic quantum mechanics, many novel quantum phenomena in systems with infinitely many degrees of freedom have been discovered [16, 17, 18, 19, 20, 21, 22].
+This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [23, 24, 25].
+These are two distinct models, referred to respectively as the tensor product algebra (TPA) model and the mutually-commuting von Neumann (observable) algebra (MCvNA) model.
+In the MCvNA model, there is, in general, no tensor product decomposition of the Hilbert space describing subsystems.
+However, it should be pointed out that relying solely on the TPA model to discuss quantum information problems has drawbacks [26, 27, 28].
+It fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom and the quantum field theory, which requires the language of type III von Neumann algebras.
+Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results from a mathematical perspective [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
+In the MCvNA model, the algebra of observables of quantum systems is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$.
+Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$ [44, 45, 19].
+It has been shown that the mutually-commuting von Neumann algebra model provides a more general framework [19].
+In the 1980s early, Summers et al.
+first introduced the maximal violation of Bell inequality and proved that its value is bounded by $2\sqrt${2} in the MCvNA model of bipartite systems, with equality attainable if and only if each algebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [46].
 This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [47].
@@
 These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [47].
-This is a pioneering work to make Bell nonlocality in quantum field theory (QFT) serve as a crucial bridge connecting quantum information science with fundamental physics [27, 33].
+This is a pioneering work to make Bell nonlocality in QFT serve as a crucial bridge connecting quantum information science with fundamental physics [47, 53].
 It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [54, 55, 56, 57, 58].
@@
 Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [59, 60, 61, 62, 63, 64].
-To date, Bell-like inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [39, 45], chain configurations [46, 47], star topologies [48, 49], polygon structures [50, 51, 52], tree-shaped networks [53, 54, 55], arbitrary acyclic networks [41, 42, 56], and arbitrary $k$-independent networks [57].
+To date, Bell-type inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [59, 65], chain configurations [66, 67], star topologies [68, 69], polygon structures [70, 71, 72], tree-shaped networks [73, 74, 75], arbitrary acyclic networks [61, 62, 76], and arbitrary $k$-independent networks [77].
 Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [78, 79].
 Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [80].
-Recently, the notion of bi-locality in an entanglement swapping network based on the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al.
-[61], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [62].
-However, Bell-like inequalities for such von Neumann algebraic models have not yet been established.
-In this paper, we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
-2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
-In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [63, 64, 65].
-This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [66, 67, 68].
-These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model.
-In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{B}}$, denoted as $\mathcal{H}=\mathcal{H}_{\mathcal{A}}\otimes\mathcal{H}_{\mathcal{B}}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators.
-Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$.
-A quantum state $\rho$ is a positive semidefinite operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$.
-The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$.
-Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
-However, in more general cases, a natural question arises: are the two models isomorphic?
-In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}$?
-This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [69, 70, 71].
-The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models.
-In 2020, Ji and collaborators [71] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative.
-This means that the tensor-product algebra model and the mutually-commuting von Neumann algebra model are not equivalent.
-Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [71, 72, 73, 74].
-3. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
-Figure 1. An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
+Recently, the notion of bilocality in an entanglement swapping network based on the MCvNA model has already been introduced by Ligthart et al.
+[81, 82], and Xu has addressed the inclusion problem between TPA model and MCvNA model in this setting [83].
+However, Bell-type inequalities in the MCvNA model have not yet been established.
+In this paper, we aim to establish bilocal inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
+2 Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
+Figure 1: An entanglement swapping network scenario with two sources. The connection between two parties represents the sharing of the physical system between them.
 Quantum bilocal scenario.
@@
 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them.
-The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively, where $x,y,z,a,b,c\in\{0,1\}$.
-Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively.
-Here it is required that the spectra of operators $A_{x},B_{y},C_{Z}$ are all {-1,1}, implying that $I\leq A_{x},B_{y},C_{z}\leq I$.
+The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively.
+Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively, with $x,y,z,a,b,c\in\{0,1\}$.
+Here it is required that the spectra of operators $A_{x},B_{y},C_{z}$ are all $\{-1,1\}$, implying that $I\leq A_{x},B_{y},C_{z}\leq I$.
 The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$.
@@
 In order to detect non-bilocal correlations generated by the network, it is often necessary to find suitable measurements that violate the following bilocal inequality
-(3.1) $\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$
+$\mathcal{S}\equiv\sqrt{|I|}+\sqrt{|J|}\leq 2,$ (2.1)
 whose maximum quantum violation is $2\sqrt{2}$ and is attainable. Here
@@
 as introduced in Ref. [63]:
-$\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
-Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurement (POVMs) performed by Alice, Bob and Charles, respectively.
-Furthermore, it follows from Ref. [57] that the violation of the inequality
-(3.2) $\mathcal{S}^{\prime}=|K|+|L|\leq 2$
-can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [78].
-Notably, different inequalities exhibit varying capabilities in detecting nonlocality.
+$\langle A_{x}B_{y}C_{z}\rangle$ $=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})$
+$=$ $\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
+Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurements (POVMs) performed by Alice, Bob and Charles, respectively.
 Mutually-commuting von Neumann algebra models.
@@
 [81, 83].
-Definition 1.
+Definition 2.1.
 (Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$.
@@
 1.
-We note that there is no correlation between the parties $\mathcal{A}$ and $\mathcal{C}$ in the network.
-Mathematically, this independence can be described by the following formula.
-So the network state of the entanglement swapping network $\tau$ should be a state in the dual space $\mathcal{M}_{\mathcal{ABC}}^{*}$, satisfying
-() $\tau(AC)=\tau(A)\tau(C)$
-for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call it the independent condition (*). This assumption will be used throughout this paper.
-Definition 2.
-A ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with the independence condition (*).
-4. Bilocal inequalities and their bounds
+We note that there is no correlation between the parties Alice and Charles in the network.
+Mathematically, this independence can be described by the following formula: a network state $\tau$ of the entanglement swapping network should be a state in $\mathcal{M}_{\mathcal{ABC}}^{*}$, the dual space of $\mathcal{M}_{\mathcal{ABC}}$, satisfying
+$\tau(AC)=\tau(A)\tau(C)$ ()
+for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call (*) the independent condition. This assumption will be used throughout this paper.
+Definition 2.2.
+The ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with all states satisfying the independence condition (*).
+3 Bilocal inequalities and their bounds
 In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network.
-Specifically, in this model, we can construct bilocal inequalities analogous to those in the non-relativistic setting.
-Define
-$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),\ J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
-where $\tau$ is the state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*).
+Specifically, in this model, we can construct the bilocal inequality analogous to that in the non-relativistic setting.
+Let
+$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),$
+$J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
+where $\tau$ is a state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*).
 Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively.
@@
 In the TMCvNA model, analogous to Ineq. (2.1), we set
-(4.1) $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$
-Analogously, corresponding to Ineq. (3.2), we define
-(4.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|,$
-where, $K_{\tau}=\tau(A_{0}(B_{0}+B_{1})C_{0}),\ L_{\tau}=\tau(A_{1}(B_{0}-B_{1})C_{1})$.
-We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$.
-The same criterion applies to $\mathcal{S}_{\tau}^{\prime}$ in Ineq.
-(4.2).
-Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively.
-The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the supremum of $\mathcal{S}_{\tau}$ defined in Ineq.
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}.$ (3.1)
+We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $\mathcal{S}_{\tau}>2$.
+The following conclusion indicates that in the TMCvNA of entanglement swapping networks, the supremum of $\mathcal{S}_{\tau}$ defined in Eq.
 (3.1) is $2\sqrt{2}$.
 This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq.
-(3.1) attains a maximal violation of $2\sqrt{2}$ based on quantum resources.
-Theorem 1. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
-Proof.
-According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{B}(\mathcal{H}_{\tau})$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$.
+(2.1) attains a maximal violation of $2\sqrt{2}$ allowed by quantum resources.
+Theorem 3.1.
+For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, we always have $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
+Proof According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{H}_{\tau}$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$.
 It follows by applying the Cauchy-Schwarz inequality that
@@
 $=2\sqrt{2}.$
-[…]
-Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Ineq.
+The final inequality invokes the condition that $I\leq A_{i}\leq I,\ -I\leq C_{i}\leq I$ and the positivity property of $\tau$. $\Box$
+Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Eq.
 (3.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$.
@@
 showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
-Theorem 2. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
-(1) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
-$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2;$
-(2) if $\mathcal{M}_{\mathcal{B}}$ is Abelian, then
-$\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|\leq 2.$
-Proof. To see (1), since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
+Theorem 3.2.
+For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
+$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
+Proof Since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
 $A_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}A_{0})(1+\epsilon_{1}A_{1}),\ \ \ C_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{1}C_{0})(1+\epsilon_{1}C_{1})$
@@
 $=2\sqrt{1}\sqrt{1}=2,$
-where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ B_{\epsilon_{0}\epsilon_{1}}$.
-To prove (2), since $\mathcal{M}_{\mathcal{B}}$ is Abelian, the four elements
-$B_{\epsilon_{0}\epsilon_{1}}\equiv\frac{1}{4}(1+\epsilon_{0}a_{0})(1+\epsilon_{1}a_{1}),$
-with $\epsilon_{0},\epsilon_{1}\in\{+,-\}$ are positive. It follows from a direct computation that
-$B_{0}+B_{1}=2(B_{++}-B_{--}),\ B_{0}-B_{1}=2(B_{+-}-B_{-+}).$
-So one obtains that
-$\mathcal{S}_{\tau}^{\prime}$ $=$ $\tau[A_{0}(B_{0}+B_{1})C_{0}]+\tau[A_{1}(B_{0}-B_{1})C_{1}]$
-$=$ $2(\tau[A_{0}(B_{++}-B_{--})C_{0}]+\tau[A_{1}(B_{+-}-B_{-+})C_{1}])$
-$\leq$ $2(\tau[B_{++}]+\tau[B_{--}]+\tau[B_{+-}]+\tau[B_{-+}])=2$
-[…]
-Remark.
-The above theorem illustrates a phenomenon: the violations of the two Bell-like inequalities, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ and $2<\mathcal{S}^{\prime}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras.
-Applying the theorem, we can infer from a violation of Ineq.
-(4.1) that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian; Similarly, a violation of Ineq.
-(4.2) implies that $\mathcal{M_{B}}$ is non-abelian.
-Here, one also needs to consider the chosen observables and the specific state.
-Nevertheless, it is important to note that these conclusions depend not only on the algebraic structure but also on the specific choice of observables and the particular state under consideration.
-To isolate the influence of the algebraic structure and examine separately the role of observables and the state in determining the values $\mathcal{S}_{\tau}$ and $\mathcal{S}_{\tau}^{\prime}$ in Ineqs.
-(4.1) and  (4.2), the following conclusion can be regarded as a generalization of the results in the non-relativistic quantum mechanical framework.
-For the sake of clarity, we introduce a tensor product structure into the TMCvNA model.
-Before presenting the following result, we first introduce the definition of separable states.
-Definition 3.
-(Separable states) Let $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ be mutually commuting von Neumann subalgebras of $\mathcal{M}_{\mathcal{AB}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}$.
-A state $\tau\in\mathcal{M}_{\mathcal{AB}}^{*}$ is said to be separable if there exist states $\xi_{\alpha}\in\mathcal{M}_{\mathcal{A}}^{*}$, $\eta_{\alpha}\in\mathcal{M}_{\mathcal{B}}^{*}$, and positive coefficients $\lambda_{\alpha}$ with $\sum_{\alpha}\lambda_{\alpha}=1$ such that for all $A\in\mathcal{M}_{\mathcal{A}},B\in\mathcal{M}_{\mathcal{B}}$,
-$\tau(AB)=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A)\eta_{\alpha}(B).$
-[…]
-$\sup_{\{A_{x},B_{y},C_{z}\}}\mathcal{S}_{\tau}=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})=2.$
-[…]
-$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
-$=\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}+A_{1})\eta_{\alpha}(B_{0}^{L})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{0}^{R})N_{\beta}(C_{0}+C_{1})|}$
-$\hskip 18.49988pt+\sqrt{|\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A_{0}-A_{1})\eta_{\alpha}(B_{1}^{R})||\sum_{\beta}\mu_{\beta}M_{\beta}(B_{1}^{R})N_{\beta}(C_{0}-C_{1})|}$
-$\leq\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}+A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}+C_{1})|)}+\sqrt{(\sum_{\alpha}\lambda_{\alpha}|\xi_{\alpha}(A_{0}-A_{1})|)(\sum_{\beta}\mu_{\beta}|N_{\beta}(C_{0}-C_{1})|)}$
-$\leq\sqrt{\sum_{\alpha}\lambda_{\alpha}(|\xi_{\alpha}(A_{0}+A_{1})|+|\xi_{\alpha}(A_{0}-A_{1})|)}\sqrt{\sum_{\beta}\mu_{\beta}(|N_{\beta}(C_{0}+C_{1})|+|N_{\beta}(C_{0}-C_{1})|)}$
-$=2\sqrt{\sum_{\alpha}\lambda_{\alpha}\max\{\xi_{\alpha}(A_{0}),\xi_{\alpha}(A_{1})\}}\sqrt{\sum_{\beta}\mu_{\beta}\max\{N_{\beta}(C_{0}),N_{\beta}(C_{1})\}}$
-$\leq 2.$
-[…]
-From the above conclusion, we see that whether a bilocal inequality can be violated depends on the choice of observables and the state, and even more so on the structure of the algebra.
+where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy-Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ C_{\epsilon_{0}\epsilon_{1}}$.
+$\Box$
+The above theorem illustrates a phenomenon: the violation of the Bell-type inequality, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras.
+Applying the theorem, we can infer from a violation of the inequality in Theorem 3.2 that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian.
 In the following, we define a quantity
-$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})$
-that appears in Theorem 3.
-Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity $\mathcal{S}^{\prime}_{\tau}$ in Ineq. (4.2).
-Corollary 4. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
+$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
+Combining Theorems 3.1 and 3.2, one naturally obtains the following corollary.
+Corollary 3.3.
+In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
 (1) for any state $\tau$ and any choice of observables $A_{x}\in\mathcal{M}_{\mathcal{A}},\ B_{y}\in\mathcal{M}_{\mathcal{B}},\ C_{z}\in\mathcal{M}_{\mathcal{C}}$ in a scheme with two inputs and two outputs, the quantity $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})$ satisfies
@@
 where $k$ is a positive constant.
-Consequently, the functional $\phi\rightarrow\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
-Proof. To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 2, we obtain (1).
-(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 2.
+Consequently, the functional $\phi\mapsto\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
+Proof To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 3.1, we obtain (1).
+(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 3.2.
 To prove (3), note that by the representations of $I_{\tau}$ and $J_{\tau}$, together with the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and the triangle inequality, one can obtain
@@
 $\leq k\sqrt{\|\phi-\psi\|},$
-[…]
-5. Maximal violation of bilocal inequalities and algebraic structures
-In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra.
-Here, violation refers to exceeding the upper bound of 2 for the $\mathcal{S}_{\tau}$ in inequality (4.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
-In the following theorem, we attempt to analyze the conditions for maximal violation.
-Theorem 5. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\tau$ is faithful, then
+where the fourth one follows the Cauchy Schwarz inequality, and the last obeys the norm for elements of $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are bounded.
+So $\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous in the state $\phi$.
+$\Box$
+4 Maximal violation of bilocal inequalities and algebraic structures
+In this section, we aim to show that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra.
+Here, violation refers to exceeding the bound for $\mathcal{S}_{\tau}$ in Eq.
+(3.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
+The following theorem analyzes the conditions for maximal violation.
+Theorem 4.1.
+For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if the network state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ is faithful, then
 $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$
 if and only if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ with $i\in\{0,1\}$.
-Remark.
-This theorem states that when the inequality is maximally violated, $A_{0},\ A_{1}$ and $\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{\mathcal{A}}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{\mathcal{A}}$.
-The same holds for $C_{0},\ C_{1}$ and $\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{M}_{\mathcal{C}}$.
-Moreover, $\tau(A_{0}A_{1}+A_{1}A_{0})=0,\ \tau(C_{0}C_{1}+C_{1}C_{0})=0.$
-Proof. If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 1 that for any $t\in[0,1]$, these equalities hold:
-(5.1) $|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$
-(5.2) $\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$
-(5.3) $\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$
-(5.4) $\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$
-(5.5) $\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$
-(5.6) $\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$
-(5.7) $A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$
+Proof If $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$, it follows from the proof of Theorem 3.1 that for any $t\in[0,1]$, these equalities hold:
+$|\tau(B_{0}(A_{0}+A_{1})(C_{0}+C_{1}))|=|\tau(B_{1}(A_{0}-A_{1})(C_{0}-C_{1}))|,$ (4.1)
+$\pi_{\tau}(B_{0})\Omega=k_{0}\pi_{\tau}[(A_{0}+A_{1})(C_{0}+C_{1})]\Omega,$ (4.2)
+$\pi_{\tau}(B_{1})\Omega=k_{1}\pi_{\tau}[(A_{0}-A_{1})(C_{0}-C_{1})]\Omega,$ (4.3)
+$\|\pi_{\tau}(B_{0})\Omega\|=\|\pi_{\tau}(B_{1})\Omega\|=1,$ (4.4)
+$\tau[(A_{0}+A_{1})^{2}]=t\tau[(C_{0}+C_{1})^{2}],$ (4.5)
+$\tau[(A_{0}-A_{1})^{2}]=t\tau[(C_{0}-C_{1})^{2}],$ (4.6)
+$A_{0}^{2}+A_{1}^{2}=2I,C_{0}^{2}+C_{1}^{2}=2I.$ (4.7)
 From (4.7), one gets $A_{i}^{2}=I,\ C_{i}^{2}=I\ (i=0,1)$ because $I\leq A_{i},C_{i}\leq I$, therefore $\tau(A_{i}^{2}A)=\tau(A),\ \tau(C_{i}^{2}C)=\tau(C)$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$.
@@
 i.e.,
-(5.8) $|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$
+$|k_{0}|=\frac{1}{2+\tau(X)},\ \ \ |k_{1}|=\frac{1}{2-\tau(X)}$ (4.8)
 because of $\tau(X)=\tau(Y)$, where $X=A_{0}A_{1}+A_{1}A_{0},\ Y=C_{0}C_{1}+C_{1}C_{0}$. Then according to Eqs. (4.2), (4.3), and (4.1),
@@
 $A_{0}A_{1}+A_{1}A_{0}=C_{0}C_{1}+C_{1}C_{0}=0,$
-and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ .
-[…]
-The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor.
-In contrast, the follwing theorem presents a highly specialized scenario in which a precise characterization of algebras attaining maximal viloation can be obtained.
-Note that next we denote by $\Lambda_{1}\overline{\otimes}\Lambda_{2}$ the von Neumann algebra of the tensor product of $\Lambda_{1},\Lambda_{2}$
-Theorem 6.
-In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if there are hyperfinite type II1 factors $\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{C}}$ such that $\mathcal{M}_{\mathcal{A}}\simeq\mathcal{M}_{\mathcal{A}}\overline{\otimes}\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}\simeq\mathcal{M}_{\mathcal{C}}\overline{\otimes}\mathcal{R}_{\mathcal{C}}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.
-Proof.
-We first deal with the special case: $\mathcal{M}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{R}_{\mathcal{C}}$.
-Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [79], $\mathcal{R}_{\mathcal{A}}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$.
-Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$.
-We can obtain that there is $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$.
-Now, from Theorem 5, we know that $A_{0},\ A_{1}$ and $A_{2}=-\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $C_{0},\ C_{1}$ and $C_{2}=-\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{K}_{\alpha}$, and $B_{0},B_{1}$ with $B_{i}^{2}=I$ such that the quantity $\mathcal{S}_{\tau}$ in Ineq.
-(4.1) arrives at $2\sqrt{2}$.
-Applying similar technologies of Eqs.
-(6)-(8) of Ref.[77], one can obtain that
-(5.9) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}+A_{1})B_{0}(C_{0}+C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2},$
-(5.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}-A_{1})B_{1}(C_{0}-C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}.$
-[…]
-$\widetilde{T}_{\alpha}=(\widetilde{A}_{0,\alpha}+\widetilde{A}_{1,\alpha})\widetilde{B}_{0,\alpha}(\widetilde{C}_{0,\alpha}+\widetilde{C}_{1,\alpha}),$
-$\widetilde{S}_{\alpha}=(\widetilde{A}_{0,\alpha}-\widetilde{A}_{1,\alpha})\widetilde{B}_{1,\alpha}(\widetilde{C}_{0,\alpha}-\widetilde{C}_{1,\alpha}).$
-Furthermore, $\Xi=\bigotimes_{\alpha\in\mathbb{N}}\Xi_{\alpha},\Upsilon=\bigotimes_{\alpha\in\mathbb{N}}\Upsilon_{\alpha},$ with $\Xi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Upsilon_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$.
-By the definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$.
-But one has
-$\sqrt{\langle\Xi,\widetilde{T}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2},\sqrt{\langle\Xi,\widetilde{S}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2}.$
-for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Xi$ and $\Upsilon$).
-[…]
-[…]
-It is mentioned that the results of Theorem 6 can be applied to quantum field theory.
-In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions, are typically type III1 factors.
-By the classical result of A.
-Connes, all injective infinite factors (except some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\overline{\otimes}\mathcal{R}_{1}$).
-Corollary 7.
-Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ be the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector.
-If $\mathcal{A}(\mathcal{O}_{1})$ and $\mathcal{A}(\mathcal{O}_{3})$ are both type III1 factors, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
-Acknowledgement. Authors thank for your comments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271394).
+and implies that $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
+It is straightforward to prove the converse process, as we check that Eqs.
+(4.1)-(4.7) hold if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$.
+We complete the proof.
+$\Box$
+To further elucidate the algebraic relations presented in Theorem 4.1, we provide the following corollary.
+Corollary 4.2.
+For the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, the bilocal inequality can be maximally violated if and only if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ contain subalgebras isomorphic to $M_{2}(\mathbb{C})$ and there exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$ that satisfies the independent condition (*): $\tau(AC)=\tau(A)\tau(C)$ for all $A\in\mathcal{M}_{\mathcal{A}},\ C\in\mathcal{M}_{\mathcal{C}}$.
+Proof Note that for any von Neumann algebra, there always exists a faithful state $\tau\in\mathcal{M}_{\mathcal{ABC}}^{*}$.
+$(\Leftarrow)$ Suppose $\mathcal{M}_{\mathcal{A}}$ (resp.
+$\mathcal{M}_{\mathcal{C}}$) contains a subalgebra $\mathcal{M}_{\mathcal{A}}^{sub}\simeq M_{2}(\mathbb{C})$ (resp.
+$\mathcal{M}_{\mathcal{C}}^{sub}\simeq M_{2}(\mathbb{C})$) and $\tau$ satisfies ()(*).
+Then there exist operators $A_{0},A_{1},$ and $A_{2}:=-\frac{{\rm{i}}}{2}[A_{0},A_{1}]$ in $\mathcal{M}_{\mathcal{A}}^{sub}$ such that they anticommute and $A_{i}^{2}=I$ for $i\in\{0,1,2\}$, where ${\rm{i}}^{2}=-1$.
+Consequently, we obtain
+$\tau(A_{i}^{2}A)=\tau(A),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$
+for any $A\in\mathcal{M}_{\mathcal{A}}$. Similarly for $C\in\mathcal{M}_{\mathcal{C}}$ with $C_{0},C_{1},C_{2}:=-\frac{{\rm{i}}}{2}[C_{0},C_{1}]$.
+Setting $B_{0}=B_{1}=I$ gives $\tau(B_{i}^{2}B)=\tau(B)$ for all $B\in\mathcal{M}_{\mathcal{B}}$.
+By Theorem 4.1, these operators yield the maximal violation $2\sqrt{2}$ of the quantity $\mathcal{S}_{\tau}$ in Eq.
+(3.1).
+$(\Rightarrow)$ Conversely, assume $\mathcal{S}_{\tau}$ in Eq. (3.1) attains the maximal violation $2\sqrt{2}$.
+Then for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ and $i\in\{0,1\}$, we have $\tau(A_{i}^{2}A)=\tau(A),\tau(B_{i}^{2}B)=\tau(B),\tau(C_{i}^{2})=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$.
+Taking $A=A_{0}A_{1}+A_{1}A_{0}$ and using the faithfulness of state $\tau$ together with $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, one gets
+$A_{0}A_{1}+A_{1}A_{0}=0,$
+i.e., $A_{0}A_{1}=-A_{1}A_{0}$. The algebra generated by $A_{0},A_{1}$ is
+$\mathfrak{A}(A_{0},A_{1}):=\{\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}|\alpha_{k}\in\mathbb{C},m,n\in\mathbb{N}\}.$
+From $\tau(A_{i}^{2}A)=\tau(A)$ and $I\leq A_{i}\leq I\ (i\in\{0,1\})$, setting $A=I$ gives $\tau(I-A_{i}^{2})=0$.
+Faithfulness of state $\tau$ then implies $A_{i}^{2}=I$.
+The same reasoning yields $B_{i}^{2}=I\ \textmd{and}\ C_{i}^{2}=I$.
+Because $A_{i}^{2}=I$, one gets $\sum_{k}\alpha_{k}A_{0}^{m}A_{1}^{n}=\alpha_{0}I+\alpha_{1}A_{0}+\alpha_{2}A_{1}+\alpha_{3}A_{0}A_{1}$. Hence,
+$\mathfrak{A}(A_{0},A_{1})={\rm span}\{I,A_{0},A_{1},-\frac{{\rm{i}}}{2}[A_{0},A_{1}]\}\simeq M_{2}(\mathbb{C}).$
+Analogously, $\mathfrak{A}(C_{0},C_{1})\simeq M_{2}(\mathbb{C})$. So this proof is completed. $\Box$
+The above results show a close connection between the maximal violation of Bell-type inequalities and algebraic structures.
+This connection can be more clearly demonstrated when all algebras are factors.
+Corollary 4.3.
+For the ternary tensor product algebra models of entanglement swapping networks, i.e., $\mathcal{M}_{\mathcal{ABC}}=\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, assume that $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ are factors.
+Then the bilocal inequality cannot be maximally violated if and only if either $\mathcal{M}_{\mathcal{A}}$ or $\mathcal{M}_{\mathcal{C}}$ is type I2k+1 for some finite positive integer $k$.
+5 Discussion and Conclusions
+We investigate Bell type inequality for entanglement swapping network in von Neumann algebraic framework, extending the paradigm that Bell violation from observable algebra structure (notably type III factors).
+We identify algebraic constraints governing inequality violation, linking network nonlocality to the noncommutative structure of the underlying algebras, and further show that maximal violation conditions can reverse-engineer the type classication of von Neumann algebras.
+This work represents merely the beginning of a much broader inquiry.
+Our current model focuses primarily on the simplest nontrivial network: the entanglement swapping scenario with two independent sources.
+The generalization of these results to arbitrary multipartite quantum networks remains an important open problem.
+In more complex architectures, the interplay between multiple independent sources and the commutation relations of their associated algebras is expected to reveal even richer structures of nonlocality in networks represented by mutually-commuting von Neumann algebras.
+Conflict of Interest  The authors declare no conflict of interest.
+Acknowledgements.
+We thank the referees for their time and comments.